Suppose the magnetic field changes uniformly from 0.500T to 0.200 T in the next 0.600s. Compute (a) the induced emf in the coil and (b) the magnitude and direction of the induced current.

1
Expert's answer
2021-04-26T17:06:25-0400

Change in Magnetic Field, "\Delta B=0.500-0.200=0.300\space T"

Time, "t=0.600\space s"


(a) Induced emf "= nA\dfrac{\Delta B}{\Delta t}=1\times\pi r^2\times\dfrac{0.300}{0.600}=\dfrac{\pi r^2}{2}"


(b) Induced current = "\dfrac{Induced \space emf}{Resistance}=\dfrac{\pi r^2}{2R}"

Since the values of r and R are not provided, therefore the direction of induced current cannot be determined.